OpenStudy (anonymous):

-4 square root of 24 -3 over 4-3square root of 24 times its conjugate

OpenStudy (valpey):

Try the equation editor.

OpenStudy (valpey):

\[-4\sqrt{24}-\frac{3}{4-3\sqrt{24}}\]times its conjugate?

OpenStudy (anonymous):

\[-4\sqrt{24}-\frac{3 }{ (4-3\sqrt{24}) (4+3\sqrt{24}) }\] @camjan is this ur problem?

OpenStudy (anonymous):

n0

OpenStudy (anonymous):

-4 square root of 24 minus 3 all over 4-3 squareroot of 24 times its conjugate which is 4+3 squareroot of 24 over 4+3 squareroot of 24

hartnn (hartnn):

is it \[\frac{-4\sqrt{24}-3}{4-3\sqrt{24}}\frac{(4+3\sqrt{24})}{(4+3\sqrt{24})}\] ?

OpenStudy (anonymous):

yes! hehehe

OpenStudy (anonymous):

pls help me answer it :)

hartnn (hartnn):

denominator can be simplified first,whats (a+b)(a-b) ?

OpenStudy (anonymous):

a squared - b squared? hehehe

hartnn (hartnn):

yes hehehe so \[(4-3\sqrt{24})(4+3\sqrt{24})=4^2-(3\sqrt{24})^2=??\]

OpenStudy (anonymous):

and then? ehehe :D

OpenStudy (anonymous):

Oh! oh! i get it already .. thankyou guys! :D

hartnn (hartnn):

glad to hear that,welcome :)